I need help with the square root property
(y+2)2-49=0

Answer by stanbon(75887)

I need help with the square root property
(y+2)^2-49=0
---------------
(y+2)^2 = 49
Take the square root of both sides to get:
y+2 = 7 or y+2 = -7
y = 5 or y = -9
=====================
Cheers,
Stan H.